OpenStudy (anonymous):

If an object moving in one direction slows down, then the direction of its acceleration is_____________________ the direction of its motion.

OpenStudy (anonymous):

NEGATIVE ACCELERATION is the result of a Force acting on an Object in the negative direction (according to whatever convention for direction we have adopted). "Does a negative acceleration mean the object is slowing down?" No, not always. First, "slowing down" is not a precise physics term; but a negative Force applied to an Object moving in a negative direction would increase the Object's speed in the negative direction. "Does a negative acceleration mean the object is moving in the opposite direction?" No. A negative Force can be applied to an Object moving in any direction.

OpenStudy (anonymous):

negative acceleration is the physics term for it, but most people sat deceleration

OpenStudy (anonymous):

*say oops

OpenStudy (karategirl2002):

cceleration can be positive while velocity is negative for an object. For example, If the object (like a car) is moving in a negative direction (let's say west) and the brakes are applied to slow the car down. This causes the car to accelerate (commonly called deceleration) in the positive direction until it stops. Mathematically, one could work this out by taking the change in velocity over the change in time. acceleration = (V final - V initial) / time In this example, let's suppose V initial is -20 m/s and V final is -10 m/s, and the time for the velocity change is 5 seconds. Then acceleration would be : -10m/s - (- 20 m/s) divided by 5 seconds = 2m/s/s Thus, we see, an example of negative velocity yielding positive acceleration from an outside net force on the object via the brakes.

OpenStudy (karategirl2002):

or NEGATIVE ACCELERATION is the result of a Force acting on an Object in the negative direction (according to whatever convention for direction we have adopted). "Does a negative acceleration mean the object is slowing down?" No, not always. First, "slowing down" is not a precise physics term; but a negative Force applied to an Object moving in a negative direction would increase the Object's speed in the negative direction. "Does a negative acceleration mean the object is moving in the opposite direction?" No. A negative Force can be applied to an Object moving in any direction. A LONGER ANSWER: --------------------------------- Do you get the feeling that there are too many different answers (among experts) for what should be an easy question? Often that means that we're not being careful about defining our terms. Let's start with "acceleration": The definition of "acceleration" is complicated by the fact that it is a commonly used English word: COMMON USAGE: acceleration: speeding up; it never means slowing down, it doesn't have any "direction" attached to it, and you'd never say, "Wow! Look at that car negatively accelerate!" PHYSICS DEFINITION: acceleration: a = Δv/Δt, or the change of Velocity per unit time, or the rate of change of Velocity with time. (The Δt here always means t_later - t_earlier and is always therefore positive ("time goes forward"); likewise, Δv means v_later - v_earlier. Here I'll use vf - vi, final minus initial.) NOTE: 1. Δv can either be positive or negative--vf can be larger, equal to, or smaller than vi. 2. Velocity is a vector--that means it has both magnitude AND direction. 3. Understanding Acceleration depends upon a good understanding of Velocity. So let's look at Velocity in terms of the defining equation for Acceleration: (1) a = Δv/Δt, where Δt > 0 always (positive). In what situations can "a" be negative? Well, that will occur whenever: (2) (vf - vi) = Δv < 0 And under what conditions will this happen? Here are some sample cases: (2a) vi, vf are both positive, and vi > vf (e.g., vi = +5 & vf = +3, so Δv = -2) (2b) vi, vf are both negative, and vi > vf (e.g., vi = -3 & vf = -5, so Δv = -2) (2c) sometimes when there's a change in direction (e.g., vi = +1 & vf = -1, so Δv = -2) SO: The term "negative acceleration" can occur in physics; in all 3 examples above, "a" is negative. The question is, What does it mean? Can we give it a simple definition? Because "negative" means "-" (unary operator), we need to explore how the signs (+ and -) are assigned to the examples in (2)? They determine the sign of Δv and therefore of "a" and therefore whether it's negative or not. And the ANSWER is simply that we NEED to assign a direction to "+" (up or down, east or west, left or right, into the page or out of it, etc.); then "-" (the negative) will mean the opposite. This convention must hold for the whole problem situation at any time--for displacement s, for velocity v, for force F, etc.). Most of the time we assume such conventions without stating them, and most of the time it works . . . simply because they are conventions. But, where we're defining a concept that is vector-related, we need to be explicit about direction. Back to Acceleration: To understand its direction better in a particular situation, I like to bring in another concept, which akm69 (another answerer) mentioned--namely Force. Early in physics we learn that: (3) F = m * a. We know that Force is a vector but Mass isn't, so therefore Acceleration must be. And, because there's no "-" in the equation (such as in F = -k * x), Acceleration and Force are in the same direction. Furthermore, any Acceleration is produced by a Force (acting in the same direction). So I like to think of "negative acceleration" in terms of Force: THE BOTTOM LINE: ---------------------- NEGATIVE ACCELERATION is the result of a Force acting on an Object in the negative direction (according to whatever convention we have adopted). In our examples (2a-c) above, let Velocity be represented on the usual x-axis with + to the right (east). The Force which produces a negative Δv (=-2) will be applied towards the left (negative) in each case. In (2a) it slows the object down: |vi| = 5 & |vf| = 3; in (2b) it speeds the object up but in the negative direction : |vi| = 3 & |vf| = 5; in (2c) it changes the object's direction by pushing toward the left: |vi| = 1 & |vf| = 1.

OpenStudy (wolfe8):